OpenStudy (yanni):

Find the slope of the line 2x+5y=-3.

OpenStudy (kropot72):

Can you put the equation in standard form? y = mx + b

OpenStudy (yanni):

5y=2x+b

OpenStudy (kropot72):

Not really. First do this: 2x + 5y = -3 Subtract 2x from each side: 2x - 2x + 5y = -3 -2x 5y = -2x -3 Now what do you get when you divide both sides bvy 5?

OpenStudy (carniel):

If you always have a hard time trying to find the slope of a line always put it in standard form.

OpenStudy (kropot72):

@yanni can you divide both sides of (5y = -2x -3) by 5?

OpenStudy (kropot72):

\[\frac{5y}{5}=?\] \[\frac{-2x-3}{5}=?\]

OpenStudy (yanni):

-2/5 -3/5

OpenStudy (yanni):

y= -2/5x - 3/5

OpenStudy (kropot72):

Excellent work! You now have the equation in standard form y = mx + b where m is the slope and b is the intercept on the y axis. Can you state the value of m now?

OpenStudy (kropot72):

The slope is the coefficient of \[-\frac{2}{5}x\]

OpenStudy (yanni):

is that the answer

OpenStudy (kropot72):

The slope is \[-\frac{2}{5}\]

OpenStudy (yanni):

so the slope of -2/5x

OpenStudy (yanni):

i mean the slope of 2x+5y=-3 is -2/5x

OpenStudy (kropot72):

The slope is -2/5 which is the coefficient of \[-\frac{2}{5}x\]